### v1

# Coherent Source Subsampling: A Data-Driven Strategy for Restoring Causal–Acausal Symmetry in Ambient Seismic Wavefield Correlations

Abstract
Ambient noise tomography relies on the assumption that the seismic wavefield is equipartitioned, meaning that energy is uniformly distributed among all directions. However, in practice, ambient noise sources are highly non-uniform in both spatial and temporal dimensions, resulting in biased estimation of the Green’s function between stations. We introduce a data-driven method, Coherent Source Subsampling (CSS), which selects and averages only those cross-correlation time windows that are associated with the excitation of sources in the stationary-zone. By confining the ensemble average to these coherent subsets, CSS effectively mitigates the influence of anisotropic or intermittent sources and restores causal–acausal symmetry in the retrieved Green’s functions. Applications to regional-scale ambient noise datasets demonstrate that CSS boosts inter-station coherence and enhances the reliability of surface-wave dispersion measurements, providing a physically interpretable bridge between source statistics and noise correlation theory.
1 Introduction
Surface-wave tomography has long been a cornerstone of the imaging of the Earth’s crust and upper mantle. The resolution of traditional surface-wave tomography, which is based on seismic waves generated by earthquakes recorded by regional and global seismic networks, is fundamentally constrained by the spatial and temporal distribution of seismicity (trampert2003global; ritzwoller1998eurasian; ritzwoller2001crustal; shapiro2002monte; ekstrom2011global; mordret2013near). Over the last 20 years, ambient seismic noise cross-correlations have supplemented surface wave tomography (shapiro2004emergence; sabra2005surface; yang2007ambient; lin2008surface; yang2008structure; yao2009analysis; ritzwoller2011ambient).
Ambient noise tomography requires cross-correlating seismic noise recorded at pairs of stations. The averaged cross-correlations are then typically inverted for the subsurface velocity structure between the stations. In this paper, $\mathbf{w}_{ij}^{(m)}$ denotes the cross-correlation vector computed between stations $i$ and $j$ over a finite time window indexed using $m$:
$\mathbf{w}_{ij}^{(m)}=\mathbf{s}^{(m)}\ast\mathcal{G}_{ij}\!\left(\mathbf{x}_{s}^{(m)}\right).$ (1)
Here, $\ast$ denotes the temporal convolution. We assume a suitable duration of ambient noise recording such that the wavefield within each time window can be approximated as being generated by an effective noise source located at $\mathbf{x}_{s}^{(m)}$ with $\mathbf{s}^{(m)}$ as its autocorrelation vector. This assumption is reasonable when the window length is short compared to the timescale over which the source distribution evolves. Our analysis shows that 30-minute windows are sufficiently short for crustal-scale ambient-noise applications. In our notation, the function $\mathcal{G}_{i}$ outputs the discrete time-domain Green’s function associated with an impulsive source at $\mathbf{x}_{s}^{(m)}$ and evaluated at the $i$th station, and the function $\mathcal{G}_{ij}$ outputs the cross-correlation between the outputs of $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. For simplicity, our discussion focuses on a single component of the seismic wavefield (e.g., vertical component for Rayleigh waves), though the framework can be readily extended to other components or multi-component analysis.
We treat the realizations $\left\{\mathbf{w}_{ij}^{(m)}\right\}_{m=1}^{M_{ij}}$ as independent samples of a random vector $\mathbf{w}_{ij}$; therefore, the empirical average
$\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]=\frac{1}{M_{ij}}\sum_{m=1}^{M_{ij}}\mathbf{w}_{ij}^{(m)}$ (2)
of $M_{ij}$ time windows provides an estimator of expectation $\mathbb{E}[\mathbf{w}_{ij}]$. Many authors have demonstrated that, in theory, the empirical average in Eq. 2 converges to the interstation Green’s function, provided the ambient wavefield is isotropic, uncorrelated, and equipartitioned (campillo2003long; wapenaar2004retrieving; wapenaar2006green; snieder2010imaging). Another important theoretical result is that, when the above conditions are satisfied, $\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]$ should exhibit causal–acausal symmetry (wapenaar2004retrieving; snieder2004extracting), representing equivalence between waves traveling in both directions between the receivers. In practice, however, this symmetry is rarely achieved (bensen2007processing; stehly2006study; yao2009analysis; tsai2009establishing) because real ambient noise sources are not isotropic (as shown in Fig. 1a) with seasonal variations in their distribution and bursts of body wave energy (stehly2007traveltime; zhan2013spurious; delaney2017passive).
We write the complete cross-correlation vector $\mathbf{w}_{ij}$ as:
[…]
where $\mathrm{concat}(\cdot)$ denotes the concatenation of two or more vectors, and $\mathrm{rev}(\cdot)$ denotes the reversal of the order of vector elements. Here, $\prescript{+}{}{\mathbf{w}}_{ij}$ and $\prescript{-}{}{\mathbf{w}}_{ij}$ denote the vectors associated with the causal and acausal branches of the cross-correlation, respectively, each arranged in order of increasing lag magnitude. The lack of symmetry between the causal and acausal branches means that $\widehat{\mathbb{E}}\left[\prescript{+}{}{\mathbf{w}}_{ij}\right]\neq\widehat{\mathbb{E}}\left[\prescript{-}{}{\mathbf{w}}_{ij}\right]$. If this is the case, then interstation tomography using either of the branches is not justified, since the assumption of equipartitioning is violated. In such cases, the empirical averages are also sensitive to the source distribution and more complicated sensitivity kernels have to the used during tomography (fichtner2015source; ermert2017ambient), which is computationally expensive.
Figure 1: Coherent source subsampling (CSS) concept. a) Ambient noise sources on the boundary, distributed non-uniformly, excite receivers at $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. The dashed line marks the source state $\lambda$=11 associated with the stationary‑phase zone. For this state, red (acausal) and cyan (causal) stars denote subsampled sources that are considered for averaging. Diamonds and a dotted line represent the second state ($\lambda$=22). CSS excludes the remaining blue-dot sources while averaging as they bias the estimation of interstation response. b) Latent‑variable autoencoder model for CSS. For each branch $b$ and pair $(i,j)$, the encoder maps input cross-correlations to three latent-variable components (orange boxes). The source-state probabilities act as data‑driven selectors that identify stationary‑phase source subsets. An example with three cross-correlation windows and two source states is depicted, although their quantities can be chosen arbitrarily.
In this paper, we present a data-driven approach, termed Coherent Source Subsampling (CSS), which subsamples the set $\left\{\mathbf{w}_{ij}^{(m)}\right\}$ based on the state of ambient noise sources. We denote the source state using a discrete variable $\lambda$, and two such source states are illustrated in Fig. 1a, using dotted and dashed lines. Our workflow for a given branch $b\in\{+,-\}$, receiver pair $(i,j)$ and source state $\lambda$ is very simple:
  • Identify windows associated with a particular source state $\lambda$, in which the cross-correlations due to several sources are coherent, to form the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$;
  • and averages only those windows to obtain
    $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]=\frac{1}{\left|\prescript{b}{\lambda}{\mathcal{C}}_{ij}\right|}\sum_{m\in\prescript{b}{\lambda}{\mathcal{C}}_{ij}}\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}.$ (4)
Here, $\left|\cdot\right|$ denotes the cardinality of a set, and note that the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is not only pair specific but also branch specific, meaning the sources associated with a particular source state differ between causal and acausal branches (see Fig. 1a). A straightforward approach to conditional averaging is to partition the data according to observable external variables, such as season, month, or time of day. For example, separate averages for winter and summer months could be computed, expecting coherent cross-correlations due to ocean activity or storms within a season (yang2008characteristics; zhan2013spurious). However, such conditioning strategies may not necessarily align with the actual coherence states in the ambient wavefield.
We adopt the representation learning framework to identify coherent source states in a data-driven manner (as shown in Fig. 1b). Specifically, we use a variational autoencoder to identify the subset of windows associated with a particular source state $\lambda$, rather than on predefined external variables. This idea is demonstrated using a synthetic experiment in the Supplementary Text S1. As an example of the source state, consider the stationary-phase zone for a given branch, where the sources coherently contribute to the inter-station response. In this paper, we label the source state for this zone as $\lambda=1$. Another source state is illustrated in Fig. 1b. We show that CSS effectively restores causal–acausal symmetry when averaging cross-correlations confined to a particular source state. Specifically, we achieve
$\widehat{\mathbb{E}}\left[\prescript{+}{}{\mathbf{w}}_{ij}\mid\lambda\right]\approx\widehat{\mathbb{E}}\left[\prescript{-}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ (5)
for a choice of source state $\lambda$ and station pair $(i,j)$. Although the causal and acausal branches are redundant in the sense that they contain the same information about the interstation Green’s function, the achievement of causal–acausal symmetry is important for two reasons. First, it validates the Green’s function approximation, which is a prerequisite for robust dispersion and travel-time measurements for tomography. Second, symmetry ensures that there is no contamination from directional source effects and other transients. To our knowledge, this is the first work that achieves the theoretically predicted high degree of causal–acausal symmetry in ambient noise cross-correlations even when the underlying source distribution is highly non-uniform and time-varying. Although CSS performs nonlinear averaging (similar to e.g., phase-weighted stacking by  schimmel1997noise), it is unique in the sense that it achieves causal–acausal symmetry.
2 Data and Methods
2.1 Noise crosscorrelations
We observed that the choice of the study region is less important to the CSS framework, which is applicable to any ambient noise dataset. However, to validate the method, we used continuous ambient seismic data from the eastern United States, a stable cratonic region characterized by low to moderate seismicity and relatively homogeneous crustal structure (liang2008ambient). In this case, low-frequency vertical-component (LHZ) displacement data from $10$ stations of the US network (Figure S11) were recorded at a sampling rate of 11\,Hz over the interval from January $2005$ through December $2007$ to analyze Rayleigh-wave propagation. The waveform data were corrected for the instrument response, detrended, demeaned, and decimated to a sampling rate of 0.50.5\,Hz. The noise is segmented into non-overlapping $30$ minute windows with a total $M_{ij}$ time windows for each pair $(i,j)$.
Before cross-correlation, each window was temporally normalized using a running absolute mean. The preprocessed traces were then cross-correlated for all available station pairs using FastXC (wang2025high), with a maximum time lag of +1000+1000 s. The cross-correlations are spectrally whitened in the 0.0090.0090.2490.249\,Hz frequency band to minimize the effects of the source signature. The resulting cross-correlations $\mathbf{w}_{ij}$ were then decomposed into their causal and acausal branches as in Equation 3. Before detailing the network architecture and subsampling procedure, we will first give a simple explanation of stationary zones, highlighting the issue to the empirical average in eq. 2, and motivating the need for the conditional average in eq. 4.
2.2 Stationary-zone source state
Ambient noise sources associated with a stationary zone are crucial, and both the causal and acausal branches each possess their own stationary zone. In the presence of an isotropic source distribution, the contribution of stationary zones dominates $\mathbb{E}\left[\mathbf{w}_{ij}\right]$, leading to causal-acausal symmetry. In stationary zones, for a given source signature $\mathbf{s}$, the cross-correlation $\mathbf{w}_{ij}$ exhibits coherence across the noise sources. We now give a simple ray-theoretical argument to show that these stationary zones correspond to regions where the travel-time difference between two stations is locally constant. The travel-time difference $\tau_{ij}=T_{j}(\mathbf{x}_{s})-T_{i}(\mathbf{x}_{s})$ represents the lag time at which a specific seismic phase appears in the cross-correlation function between stations $i$ and $j$. Here, for simplicity, we ignore the frequency dependence of the lag time, which corresponds to the differential travel time of a phase propagating from the source location $\mathbf{x}_{s}$ to the two receivers.
For branch $b$, the cross-correlation vector $\prescript{b}{}{\mathbf{w}}_{ij}$ depends on two random variables: the signature of the source $\mathbf{s}$ and the location of the source $\mathbf{x}_{s}$, whose joint distribution is denoted by $\rho(\mathbf{s},\mathbf{x}_{s})$. The distribution $\rho(\mathbf{s},\mathbf{x}_{s})$ encapsulates the statistical properties of the ambient noise field, including source anisotropy and temporal variability. The expectation of the random vector $\prescript{b}{}{\mathbf{w}}_{ij}$ is given by
$\mathbb{E}\big[\prescript{b}{}{\mathbf{w}}_{ij}\big]=\iint\prescript{b}{}{\mathbf{w}}_{ij}\,\rho(\mathbf{x}_{s},\mathbf{s})\,\mathrm{d}\mathbf{x}_{s}\,\mathrm{d}\mathbf{s}=\iint\mathbb{E}\big[\prescript{b}{}{\mathbf{w}}_{ij}\mid\tau_{ij}\big]\,\rho(\tau_{ij},\mathbf{s})\,\mathrm{d}\tau_{ij}\,\mathrm{d}\mathbf{s},$ (6)
where $\mathbb{E}\big[\prescript{b}{}{\mathbf{w}}_{ij}\mid\tau_{ij}\big]$ is the conditional expectation of $\prescript{b}{}{\mathbf{w}}_{ij}$ given the difference in travel-time $\tau_{ij}$. For a specific seismic phase arrival, we can transform the source distribution using the lag time $\tau_{ij}=T_{j}(\mathbf{x}_{s})-T_{i}(\mathbf{x}_{s})$. The pushforward distribution $\rho(\tau_{ij},\mathbf{s})$ is obtained from the joint distribution $\rho(\mathbf{x}_{s},\mathbf{s})$ as:
$\rho(\tau_{ij},\mathbf{s})=\int\delta\big(\tau_{ij}-T_{ij}(\mathbf{x}_{s})\big)\,\rho(\mathbf{x}_{s},\mathbf{s})\,\mathrm{d}\mathbf{x}_{s}=\int_{T_{ij}(\mathbf{x}_{s})=\tau_{ij}}\frac{\rho(\mathbf{x}_{s},\mathbf{s})}{\|\nabla T_{ij}(\mathbf{x}_{s})\|}\,\mathrm{d}S(\mathbf{x}_{s}).$ (7)
Here, integration is performed over a manifold, where $\mathrm{d}S(\mathbf{x}_{s})$ denotes the surface element on the manifold $T_{ij}(\mathbf{x}_{s})=\mathrm{constant}$. The gradient $\nabla T_{ij}(\mathbf{x}_{s})$ measures how sensitively the travel-time delay between receivers $i$ and $j$ changes with respect to the source position $\mathbf{x}_{s}$. In stationary zones, small perturbations in the source position do not affect the delay to the first order, which implies that the $\ell_{2}$ norm of the gradient $\|\nabla T_{ij}(\mathbf{x}_{s})\|$ is close to zero. Due to this, the cross-correlations are coherent across sources located in stationary zones, leading to constructive interference. We will later leverage this coherence in the design of our neural network architecture.
Although the cross-correlations are coherent within stationary zones, in practice, the energy contribution from such zones does not necessarily dominate the ensemble average over all available windows (Eq. 2) — this is because the source density $\rho(\mathbf{s},\mathbf{x}_{s})$ may be negligibly small in those regions. In other words, as Eq. (7) indicates that the resulting distribution $\rho(\tau_{ij},\mathbf{s})$ depends jointly on the mapping $T_{ij}(\mathbf{x}_{s})$, and on the spatial source distribution $\rho(\mathbf{x}_{s},\mathbf{s})$, the departure of $\rho(\mathbf{x}_{s},\mathbf{s})$ from theoretical assumptions biases the interstation Green’s function retrieval. In such cases, the retrieval of Green’s functions can be improved by isolating and averaging only windows from stationary zones. We fix $\lambda=1$ to denote the source state of the stationary zone, which is the only state of interest in this study. Additional source states may exist, namely alternative configurations of the ambient noise field, defined by sources with similar signatures that yield coherent cross-correlations between receiver pairs for a given arrival; we label these by $\lambda=2,\ldots,K$. Fig. 1a and the synthetic experiment (Supplementary Text S11) provide representative examples of such states.
2.3 Latent variable modeling
To enable data-driven identification of coherent source subsets $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$, we adopt a latent-variable modeling approach based on variational autoencoders (VAEs). VAEs provide a probabilistic framework for learning interpretable latent representations (kingma2013auto; doersch2016tutorial; bishop2023deep; prince2023understanding) from high-dimensional data. Within this framework, we introduce a discrete latent variable $\lambda$ (morton2021scalable) to represent source states and design the network architecture such that $\lambda$ correlates with physically meaningful source states, such as the stationary-zone state. The design is based on deliberate architectural decisions: a permutation-invariant encoder that accumulates coherent information (bharadwaj2024extracting) across all windows for a given pair and a categorical posterior distribution over $\lambda$ that naturally groups coherent cross-correlation windows. We emphasize that alternative network designs may exist; however, the focus of this paper is not to exhaustively explore all possible architectures, but rather to demonstrate that a carefully designed VAE can successfully solve the CSS problem by learning to distinguish stationary-zone windows from incoherent noise. The following subsections brief the generative and inference models that implement this approach.
2.3.1 Generative model
We begin with a generative model that mathematically describes how observed cross-correlations are synthesized from latent variables. That is to say, we train a decoder network that generates each cross-correlation waveform $\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}$ using three distinct latent-variable components — they are highlighted in Fig. 1b.
  • Coherent source features $\left(\prescript{b}{\lambda}{\mathbf{g}}_{ij}\right)$. This component represents the coherent contributions associated with a particular source state $\lambda$, branch $b$ and a station pair $(i,j)$. Coherency means that this component is invariant to the ordering of time windows associated with a specific pair and branch. This invariance is achieved by inferring $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$ from the entire ensemble of $M$ cross-correlation windows for the station pair $(i,j)$ and the branch $b$.
[…]
    $\sum_{\lambda=1}^{K}\prescript{b}{\lambda}{\pi}_{ij}^{(m)}=1,$ (8)
    and are used later to identify the windows associated with each source state.
  • Nuisance features $\left(\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}\right)$. This component is specific to each time window $m$ and captures features that are not linked with any of the source states — for example, the cross-correlations due to incoherent sources away from the stationary zones.
In summary, our framework generates each cross-correlation waveform as:
$\prescript{b}{}{\widehat{\mathbf{w}}}_{ij}^{(m)}=\mathtt{decoder}\left(\prescript{b}{l}{\mathbf{g}}_{ij},\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}\right),\quad\text{where}\;l=\arg\max_{\lambda}\prescript{b}{\lambda}{\pi}_{ij}^{(m)},\quad m=1,\ldots,M_{ij}.$ (9)
Here, $l$ denotes the source state with the highest posterior probability for the window $m$, effectively selecting a single coherent feature $\prescript{b}{l}{\mathbf{g}}_{ij}$ for generation. We implement the $\mathtt{decoder}$ as a convolutional neural network.
2.3.2 Inference model
The generative model in Equation 9 defines the forward process to synthesize cross-correlations from latent variables. To complete the VAE framework, we now detail the corresponding inference model that inverts this process by estimating the latent-variable components from observed data. We perform probabilistic inference on the latent-variable components listed earlier using three components of the $\mathtt{encoder}$ network:
[…]
$\prescript{b}{}{\mathbf{n}}_{ij}^{(m)}$ $=\mathtt{enc}_{\mathbf{n}}\left(\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}\right),\quad\text{for each}\;m=1,\ldots,M_{ij}.$
Here, $\mathtt{enc}_{\mathbf{g}}$, $\mathtt{enc}_{\pi}$, and $\mathtt{enc}_{\mathbf{n}}$ are convolutional encoder components that infer coherent source features, source state probabilities, and nuisance features, respectively. The symmetric encoder component $\mathtt{enc}_{\mathbf{g}}$ exploits the permutation invariance by accumulating information across all time windows for a given station pair and branch. This component infers $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$, which is constrained to be invariant to window ordering. On the contrary, $\mathtt{enc}_{\pi}$ and $\mathtt{enc}_{\mathbf{n}}$ operate on individual windows to capture window-specific features, inferring $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ and $\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}$, respectively. Note that the autoencoder may assign the stationary-zone state to any index in $\{1,\cdots,K\}$. After training, we therefore relabeled the source states so that the stationary-zone state is always set to $\lambda=1$. Additional architectural and training details are given in the Supplementary Text S22.
2.4 Coherent source subsampling (CSS)
We employ an outer training loop over station pairs: for each pair $(i,j)$ we initialize the VAE from the previous section and train it jointly on both branches of that specific pair. After training, CSS aims to approximate the conditional expectation of each branch $b$ for pair $(i,j)$:
$\mathbb{E}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]=\iint\prescript{b}{}{\mathbf{w}}_{ij}(\mathbf{x}_{s},\mathbf{s})\,\rho(\mathbf{x}_{s},\mathbf{s}\mid\lambda)\,\mathrm{d}\mathbf{x}_{s}\,\mathrm{d}\mathbf{s},$ (11)
where the condition isolates contributions from sources associated with the state indexed by $\lambda$. It uses the learned source state probabilities $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ for subsampling. Specifically, for each branch $b$, station pair $(i,j)$, and source state $\lambda$, the CSS estimator of the conditional expectation in Eq. 11 is given by Equation 4, where the subset of correlation windows $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is defined as:
$\prescript{b}{\lambda}{\mathcal{C}}_{ij}=\left\{m:\prescript{b}{\lambda}{\pi}_{ij}^{(m)}>\alpha\right\}.$ (12)
It is important to note that the identification of windows belonging to a specific source state is accomplished by probability thresholding (we used $\alpha=0.85$). By averaging only the time windows with high posterior probability $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$, we in effect subsample the windows associated with a given source state. With this targeted subsampling and by setting $\lambda=1$, we can isolate the contributions from stationary regions in the source field. Note that the conditional average $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ in Eq. 4 is a non‑linear function of the data, unlike linear averaging, because the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ itself depends on the data. It is important to emphasize that the averaging is performed on the original cross-correlation windows, using the VAE only to identify which windows belong to each source state. Apart from the source state probabilities, we do not use any other outputs of the encoder or decoder networks, such as the reconstructed cross-correlations, in our analysis. We now demonstrate the advantages of subsampling over indiscriminate averaging.
3 Results
Figure 2: Cross-correlation gathers for the eastern United States. (a) Before subsampling: The linearly averaged cross-correlations, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, show large amplitudes close to zero lag and pronounced causal-acausal asymmetry, indicating a directionally biased ambient noise field. (b) After subsampling: The conditional averages, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$, obtained after subsampling, display enhanced waveform coherence and substantially improved causal–acausal symmetry. The percentages of time-window allocation (in red) indicate the fraction of windows selected for the source state $\lambda\text{ = 1}$. The correlation values (in blue) measure the degree of symmetry between the acausal and causal components for each station pair. The black lines mark the time lags corresponding to velocities of $2$ and 55\,km/s.
[…]
The improvement in cross-correlation symmetry translates directly into enhanced Rayleigh-wave dispersion measurements. Figures 3c–e for (GOGA, NCB) and 3k–m for (BLA, PKME) display the Rayleigh-wave dispersion images obtained without subsampling, which show asymmetry between the causal and acausal branches. After applying CSS averaging for $\lambda=1$, the dispersion images (Figures 3f–h and 3n–p) exhibit markedly enhanced coherence and symmetry. This improved balance between the causal and acausal branches enables a much more reliable picking of group velocities. Figures 4a–i show group-velocity dispersion curves, extracted by automated picking, for nine station pairs with short, intermediate, and long interstation distances. For each pair, the dispersion curves obtained by simple linear averaging display substantial scatter and a clear mismatch between the causal and acausal branches. After applying conditional averaging, the dispersion estimates become much more stable: the causal and acausal branches closely coincide, producing nearly symmetric behavior in the 5505–50\,s period band, where Rayleigh-wave energy is strongest. The pairs with the longest and intermediate distances, including (GOGA, NCB) (14201420\,km), (BLA, PKME) (12901290\,km), (CBN, PKME) (10321032\,km), and (ACSO, CNNC) (713713\,km), show the most pronounced improvements, with the subsampled curves collapsing onto a well-defined trend. Other station pairs, such as (AAM, GOGA) (989989\,km), (BINY, GLMI) (755755\,km), (AAM, BINY) (632632\,km), (ACSO, BINY) (625625\,km), and (BLA, GOGA) (505505\,km), also exhibit notably closer agreement between the causal and acausal branches than in the linear averages. Finally, we compare the CSS-derived dispersion curves with the earthquake-derived dispersion curves for (GOGA, NCB) and (ACSO, CNNC) pairs, using earthquakes of about $M_{w}\,4.5$ located near the GOGA and ACSO stations, respectively. For both pairs, the subsampled symmetric dispersion curves agree well with the earthquake-based dispersion curves.
It is important to note that while CSS enhances time-domain symmetry, it does not guarantee identical causal and acausal dispersion curves across all frequency bands. This is an expected consequence: if the stationary zone lacks sources at specific frequencies, or if certain frequency bands are naturally weak or absent in the ambient field, CSS can enhance the coherence of existing energy but cannot artificially generate it. Thus, the method optimizes the extraction of information from the available ambient field but is fundamentally constrained by the spectral content of that field.
Figure 3: Temporal and dispersion analysis for the station pairs (GOGA, NCB) and (BLA, PKME). (a–b, i–j) Acausal and causal time-window selections obtained from CSS for source states $\lambda\text{ = \{1, 2\} }$. (c–e, k–m) Before subsampling: standard linear averaging produces dispersion images and cross-correlations with low coherence and pronounced causal–acausal asymmetry. (f-h, n-p) After subsampling: source state $\lambda\text{ = 1}$ results in dispersion images and cross-correlations with substantially improved causal–acausal symmetry.
Figure 4: Group‐velocity dispersion curves for nine station pairs in the eastern United States. Solid lines represent the causal and acausal dispersion curves obtained for source state $\lambda\text{ = 1}$, while dotted lines indicate the corresponding dispersion curves based on linear averaging. Earthquake‐based dispersion curves are shown for comparison for station pairs (GOGA, NCB) and (ACSO, CNNC). For all station pairs, the dispersion curves retrieved using source state $\lambda\text{=1}$ show markedly reduced scatter between 1010\, and 5050\,s, whereas the linearly averaged curves display substantial variability and systematically underestimate velocities. This closer agreement demonstrates the effectiveness of CSS for recovering symmetric dispersion information.
4 Discussion
The results show that CSS effectively isolates the relatively small fraction of ambient-noise windows that exert the strongest control on empirical Green’s function retrieval, yielding substantial improvements in causal–acausal symmetry and dispersion stability. By extracting quasi-stable noise regimes that recur over long periods, CSS provides a physically interpretable connection between ambient-noise variability and the theoretical conditions required for reciprocal Green’s function estimation.
Because it reduces the number of windows needed to achieve coherent stacking, CSS is particularly well suited to large-scale experiments, including dense nodal arrays and distributed acoustic sensing (DAS). In such data-heavy environments, CSS can retrieve coherent surface-wave energy even in the presence of intense anthropogenic noise and pronounced spatial heterogeneity, enhancing interferometric quality while greatly decreasing data volume.
More broadly, the conditional-averaging framework at the core of CSS may benefit applications where coherent wavefield reconstruction is constrained by uncertain or varying source distributions. A current practical limitation is that our implementation trains a separate VAE for each station pair; while this is feasible for modest-sized networks, scaling to hundreds of pairs would require substantial computational effort or new joint, multi-pair training strategies. Developing such scalable approaches is a crucial step toward deploying CSS at network-wide or continental scales. In future work, CSS could also be adapted to independently subsample narrow frequency bands in order to detect frequency-specific weak sources within the stationary zone.
5 Conclusion
We introduced Coherent Source Subsampling (CSS), a data-driven approach that reformulates ambient noise averaging as a conditional process guided by source-state coherence. Instead of relying on indiscriminate stacking, CSS computes conditional averages over subsets representing quasi-stationary segments of the ambient field. By isolating these coherent intervals, CSS suppresses the influence of transient and directionally biased noise and substantially enhances the causal–acausal symmetry of the retrieved cross-correlations.
Application to the eastern United States dataset demonstrates that coherent Green-function estimates can be obtained from only a small fraction of all available time windows. Conditional averages produced through CSS yield clearer Rayleigh-wave arrivals, improved causal–acausal agreement, and more stable group-velocity dispersion curves across a range of interstation distances. These improvements reduce the bias and variability inherent in linear stacking and provide dispersion measurements better suited for robust ambient-noise tomography. In general, CSS offers a physically interpretable and empirically validated pathway for extracting reliable structural information from nonstationary and anisotropic noise fields, extending the applicability of ambient noise imaging in challenging environments.
Open Research Section
Waveform data used in this study can be obtained from the IRIS EarthScope Data Management Center (https://ds.iris.edu/ds/nodes/dmc/) through the FDSN dataselect service (http://service.iris.edu/fdsnws/dataselect/1/). The dataset includes recordings from the United States National Seismic Network (network code US) USArray. All codes developed as part of this work will be made publicly available after the manuscript undergoes peer review.
Conflict of Interest
The authors declare that there are no conflicts of interest for this manuscript.
Acknowledgments
This work was supported by the Shell project under project titled Neural Network Based Denoising Of Passive Seismic Data. We would like to thank Dr. Anu Chandran (Shell), Dr. Rene-Edouard Plessix (Shell), Dr. Arjun Datta (IISER Pune), Prof. Shyam Rai (IISER Pune), Prof. Aurélien Mordret (GEUS), Prof. Gerard Schuster (University of Utah) and Dr. Kees Weemstra (KNMI) for their valuable feedback. We thank Tiente Rengneichuong Koireng (IISc) and Pushkar Bharadwaj (IISc) for testing the method on other datasets. The authors used AI-assisted tools (OpenAI ChatGPT and Writefull) for language refinement. All edits were reviewed, and the authors are solely responsible for the final text.

### v2

@@
 Abstract
-Ambient noise tomography relies on the assumption that the seismic wavefield is equipartitioned, meaning that energy is uniformly distributed among all directions.
-However, in practice, ambient noise sources are highly non-uniform in both spatial and temporal dimensions, resulting in biased estimation of the Green’s function between stations.
-We introduce a data-driven method, Coherent Source Subsampling (CSS), which selects and averages only those cross-correlation time windows that are associated with the excitation of sources in the stationary-zone.
-By confining the ensemble average to these coherent subsets, CSS effectively mitigates the influence of anisotropic or intermittent sources and restores causal–acausal symmetry in the retrieved Green’s functions.
-Applications to regional-scale ambient noise datasets demonstrate that CSS boosts inter-station coherence and enhances the reliability of surface-wave dispersion measurements, providing a physically interpretable bridge between source statistics and noise correlation theory.
+Ambient noise tomography relies on the assumption that the seismic wavefield is equipartitioned.
+In practice, ambient noise sources are spatially and temporally heterogeneous, producing biased estimates of Green’s function between stations.
+We introduce a data-driven method, Coherent Source Subsampling (CSS), which selects and averages only cross-correlation time windows associated with excitation of sources in the stationary zone.
+By restricting the ensemble average to these windows, CSS mitigates effects of non-uniform source distribution and restores causal–acausal symmetry in the retrieved interstation response.
+Applications to regional ambient-noise datasets show that CSS stabilizes surface-wave dispersion measurements even when source statistics violate assumptions of standard seismic interferometry.
+For the central California dataset, CSS-derived group-velocity tomograms consistently image a high-velocity block between the Rinconada and San Andreas faults across multiple periods.
+In comparison, the full-ensemble (linear) average does not capture this block, which is well established.
+Our approach is particularly useful for short-duration passive surveys.
 1 Introduction
 Surface-wave tomography has long been a cornerstone of the imaging of the Earth’s crust and upper mantle.
-The resolution of traditional surface-wave tomography, which is based on seismic waves generated by earthquakes recorded by regional and global seismic networks, is fundamentally constrained by the spatial and temporal distribution of seismicity (trampert2003global; ritzwoller1998eurasian; ritzwoller2001crustal; shapiro2002monte; ekstrom2011global; mordret2013near).
-Over the last 20 years, ambient seismic noise cross-correlations have supplemented surface wave tomography (shapiro2004emergence; sabra2005surface; yang2007ambient; lin2008surface; yang2008structure; yao2009analysis; ritzwoller2011ambient).
+The resolution of traditional surface-wave tomography, which is based on seismic waves generated by earthquakes recorded by regional and global seismic networks, is fundamentally constrained by the spatial and temporal distribution of seismicity (Trampert and Woodhouse, 2003; Ritzwoller and Levshin, 1998; Ritzwoller et al., 2001; Shapiro and Ritzwoller, 2002; Ekström, 2011; Mordret et al., 2013).
+Over the last 20 years, ambient seismic noise cross-correlations have supplemented surface wave tomography (Shapiro and Campillo, 2004; Sabra et al., 2005; Yang et al., 2007; Lin et al., 2008; Yang et al., 2008; Yao and Van Der Hilst, 2009; Ritzwoller et al., 2011).
 Ambient noise tomography requires cross-correlating seismic noise recorded at pairs of stations.
@@
 of $M_{ij}$ time windows provides an estimator of expectation $\mathbb{E}[\mathbf{w}_{ij}]$.
-Many authors have demonstrated that, in theory, the empirical average in Eq.
-2 converges to the interstation Green’s function, provided the ambient wavefield is isotropic, uncorrelated, and equipartitioned (campillo2003long; wapenaar2004retrieving; wapenaar2006green; snieder2010imaging).
-Another important theoretical result is that, when the above conditions are satisfied, $\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]$ should exhibit causal–acausal symmetry (wapenaar2004retrieving; snieder2004extracting), representing equivalence between waves traveling in both directions between the receivers.
-In practice, however, this symmetry is rarely achieved (bensen2007processing; stehly2006study; yao2009analysis; tsai2009establishing) because real ambient noise sources are not isotropic (as shown in Fig.
-1a) with seasonal variations in their distribution and bursts of body wave energy (stehly2007traveltime; zhan2013spurious; delaney2017passive).
+Many authors have demonstrated that, in theory, the empirical average in Equation 2 converges to the interstation Green’s function, provided the ambient wavefield is isotropic, uncorrelated, and equipartitioned (Campillo and Paul, 2003; Wapenaar, 2004; Wapenaar and Fokkema, 2006; Snieder and Wapenaar, 2010).
+Another important theoretical result is that, when the above conditions are satisfied, $\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]$ should exhibit causal–acausal symmetry (Wapenaar, 2004; Snieder, 2004), representing equivalence between waves traveling in both directions between the receivers.
+In practice, however, this symmetry is rarely achieved (Bensen et al., 2007; Stehly et al., 2006; Yao and Van Der Hilst, 2009; Tsai, 2009) because real ambient noise sources are not isotropic (as shown in Figure 1a) with seasonal variations in their distribution and bursts of body wave energy (Stehly et al., 2007; Zhan et al., 2013; Delaney et al., 2017).
 We write the complete cross-correlation vector $\mathbf{w}_{ij}$ as:
@@
 If this is the case, then interstation tomography using either of the branches is not justified, since the assumption of equipartitioning is violated.
-In such cases, the empirical averages are also sensitive to the source distribution and more complicated sensitivity kernels have to the used during tomography (fichtner2015source; ermert2017ambient), which is computationally expensive.
+In such cases, the empirical averages are also sensitive to the source distribution and more complicated sensitivity kernels have to the used during tomography (Fichtner, 2015; Ermert et al., 2017), which is computationally expensive.
 Figure 1: Coherent source subsampling (CSS) concept.
@@
 In this paper, we present a data-driven approach, termed Coherent Source Subsampling (CSS), which subsamples the set $\left\{\mathbf{w}_{ij}^{(m)}\right\}$ based on the state of ambient noise sources.
-We denote the source state using a discrete variable $\lambda$, and two such source states are illustrated in Fig.
-1a, using dotted and dashed lines.
+We denote the source state using a discrete variable $\lambda$, and two such source states are illustrated in Figure 1a, using dotted and dashed lines.
 Our workflow for a given branch $b\in\{+,-\}$, receiver pair $(i,j)$ and source state $\lambda$ is very simple:
@@
     $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]=\frac{1}{\left|\prescript{b}{\lambda}{\mathcal{C}}_{ij}\right|}\sum_{m\in\prescript{b}{\lambda}{\mathcal{C}}_{ij}}\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}.$ (4)
-Here, $\left|\cdot\right|$ denotes the cardinality of a set, and note that the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is not only pair specific but also branch specific, meaning the sources associated with a particular source state differ between causal and acausal branches (see Fig.
-1a).
+Here, $\left|\cdot\right|$ denotes the cardinality of a set, and note that the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is not only pair specific but also branch specific, meaning the sources associated with a particular source state differ between causal and acausal branches (see Figure 1a).
 A straightforward approach to conditional averaging is to partition the data according to observable external variables, such as season, month, or time of day.
-For example, separate averages for winter and summer months could be computed, expecting coherent cross-correlations due to ocean activity or storms within a season (yang2008characteristics; zhan2013spurious).
+For example, separate averages for winter and summer months could be computed, expecting coherent cross-correlations due to ocean activity or storms within a season (Yang and Ritzwoller, 2008; Zhan et al., 2013).
 However, such conditioning strategies may not necessarily align with the actual coherence states in the ambient wavefield.
-We adopt the representation learning framework to identify coherent source states in a data-driven manner (as shown in Fig.
-1b).
+We adopt the representation learning framework to identify coherent source states in a data-driven manner (as shown in Figure 1b).
 Specifically, we use a variational autoencoder to identify the subset of windows associated with a particular source state $\lambda$, rather than on predefined external variables.
-This idea is demonstrated using a synthetic experiment in the Supplementary Text S1.
+This idea is demonstrated using synthetic experiments for non-uniform source distributions with varying noise levels in the Supplementary Text S1.
 As an example of the source state, consider the stationary-phase zone for a given branch, where the sources coherently contribute to the inter-station response.
 In this paper, we label the source state for this zone as $\lambda=1$.
-Another source state is illustrated in Fig.
-1b.
+Another source state is illustrated in Figure 1b.
 We show that CSS effectively restores causal–acausal symmetry when averaging cross-correlations confined to a particular source state.
@@
 To our knowledge, this is the first work that achieves the theoretically predicted high degree of causal–acausal symmetry in ambient noise cross-correlations even when the underlying source distribution is highly non-uniform and time-varying.
-Although CSS performs nonlinear averaging (similar to e.g., phase-weighted stacking by  schimmel1997noise), it is unique in the sense that it achieves causal–acausal symmetry.
+We further show that the improved symmetry leads to more accurate surface-wave tomographic maps.
+Finally, note that while CSS performs nonlinear averaging, similar to phase-weighted stacking by  Schimmel and Paulssen (1997), it is unique in the sense that it achieves causal–acausal symmetry.
 2 Data and Methods
 2.1 Noise crosscorrelations
-We observed that the choice of the study region is less important to the CSS framework, which is applicable to any ambient noise dataset.
-However, to validate the method, we used continuous ambient seismic data from the eastern United States, a stable cratonic region characterized by low to moderate seismicity and relatively homogeneous crustal structure (liang2008ambient).
-In this case, low-frequency vertical-component (LHZ) displacement data from $10$ stations of the US network (Figure S11) were recorded at a sampling rate of 11\,Hz over the interval from January $2005$ through December $2007$ to analyze Rayleigh-wave propagation.
-The waveform data were corrected for the instrument response, detrended, demeaned, and decimated to a sampling rate of 0.50.5\,Hz.
-The noise is segmented into non-overlapping $30$ minute windows with a total $M_{ij}$ time windows for each pair $(i,j)$.
-Before cross-correlation, each window was temporally normalized using a running absolute mean.
-The preprocessed traces were then cross-correlated for all available station pairs using FastXC (wang2025high), with a maximum time lag of +1000+1000 s.
-The cross-correlations are spectrally whitened in the 0.0090.0090.2490.249\,Hz frequency band to minimize the effects of the source signature.
-The resulting cross-correlations $\mathbf{w}_{ij}$ were then decomposed into their causal and acausal branches as in Equation 3.
-Before detailing the network architecture and subsampling procedure, we will first give a simple explanation of stationary zones, highlighting the issue to the empirical average in eq.
-2, and motivating the need for the conditional average in eq.
-4.
+We observed that the choice of the study region is less important to the CSS framework, which applies to any ambient noise dataset.
+Here, we chose to demonstrate the method by analyzing continuous ambient seismic data recorded in central California using $20$ stations from the Central California Seismic Experiment (CCSE) Transportable Array deployment (CCSE, 2015).
+This area poses a stringent challenge for the method because the ambient-noise wavefield is highly heterogeneous and strongly directional.
+The Pacific Ocean acts as a dominant and persistent source of microseismic energy, leading to a marked causal–acausal asymmetry in interstation correlations (Retailleau and Beroza, 2021).
+In conventional ambient-noise tomography, the correlation branch associated with waves arriving from the Pacific is often preferentially used for dispersion measurements because the opposite branch is weakly illuminated (Jiang et al., 2018).
+The selected stations form a west–east transect that extends from the Southern Coast Ranges (SCR), through the Rinconada (RF) and San Andreas (SAF) faults system, and into the Great Valley Basin (GVB) (Supplementary Figure S55).
+Data were processed using standard ambient-noise interferometry procedures discussed in Bensen et al.
+(2007).
+The detailed preprocessing steps are described in the Supplementary Text S22.
+Before detailing the network architecture and subsampling procedure, we first provide a simple explanation of stationary zones, highlighting the limitation of the linear empirical average in Equation 2 and motivating the conditional average in Equation 4.
 2.2 Stationary-zone source state
@@
 We will later leverage this coherence in the design of our neural network architecture.
-Although the cross-correlations are coherent within stationary zones, in practice, the energy contribution from such zones does not necessarily dominate the ensemble average over all available windows (Eq.
-2) — this is because the source density $\rho(\mathbf{s},\mathbf{x}_{s})$ may be negligibly small in those regions.
-In other words, as Eq.
-(7) indicates that the resulting distribution $\rho(\tau_{ij},\mathbf{s})$ depends jointly on the mapping $T_{ij}(\mathbf{x}_{s})$, and on the spatial source distribution $\rho(\mathbf{x}_{s},\mathbf{s})$, the departure of $\rho(\mathbf{x}_{s},\mathbf{s})$ from theoretical assumptions biases the interstation Green’s function retrieval.
+Although the cross-correlations are coherent within stationary zones, in practice, the energy contribution from such zones does not necessarily dominate the ensemble average over all available windows (Equation 2) — this is because the source density $\rho(\mathbf{s},\mathbf{x}_{s})$ may be negligibly small in those regions.
+In other words, as Equation 7 indicates that the resulting distribution $\rho(\tau_{ij},\mathbf{s})$ depends jointly on the mapping $T_{ij}(\mathbf{x}_{s})$, and on the spatial source distribution $\rho(\mathbf{x}_{s},\mathbf{s})$, the departure of $\rho(\mathbf{x}_{s},\mathbf{s})$ from theoretical assumptions biases the interstation Green’s function retrieval.
 In such cases, the retrieval of Green’s functions can be improved by isolating and averaging only windows from stationary zones.
@@
 Additional source states may exist, namely alternative configurations of the ambient noise field, defined by sources with similar signatures that yield coherent cross-correlations between receiver pairs for a given arrival; we label these by $\lambda=2,\ldots,K$.
-Fig.
-1a and the synthetic experiment (Supplementary Text S11) provide representative examples of such states.
+Figure 1a and the synthetic experiments (Supplementary Text S11) provide representative examples of such states.
 2.3 Latent variable modeling
 To enable data-driven identification of coherent source subsets $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$, we adopt a latent-variable modeling approach based on variational autoencoders (VAEs).
-VAEs provide a probabilistic framework for learning interpretable latent representations (kingma2013auto; doersch2016tutorial; bishop2023deep; prince2023understanding) from high-dimensional data.
-Within this framework, we introduce a discrete latent variable $\lambda$ (morton2021scalable) to represent source states and design the network architecture such that $\lambda$ correlates with physically meaningful source states, such as the stationary-zone state.
-The design is based on deliberate architectural decisions: a permutation-invariant encoder that accumulates coherent information (bharadwaj2024extracting) across all windows for a given pair and a categorical posterior distribution over $\lambda$ that naturally groups coherent cross-correlation windows.
+VAEs provide a probabilistic framework for learning interpretable latent representations (Kingma and Welling, 2013; Doersch, 2016; Bishop and Bishop, 2023; Prince, 2023) from high-dimensional data.
+Within this framework, we introduce a discrete latent variable $\lambda$ (Morton et al., 2021) to represent source states and design the network architecture such that $\lambda$ correlates with physically meaningful source states, such as the stationary-zone state.
+The design is based on deliberate architectural decisions: a permutation-invariant encoder that accumulates coherent information (Bharadwaj, 2024) across all windows for a given pair and a categorical posterior distribution over $\lambda$ that naturally groups coherent cross-correlation windows.
 We emphasize that alternative network designs may exist; however, the focus of this paper is not to exhaustively explore all possible architectures, but rather to demonstrate that a carefully designed VAE can successfully solve the CSS problem by learning to distinguish stationary-zone windows from incoherent noise.
@@
 We begin with a generative model that mathematically describes how observed cross-correlations are synthesized from latent variables.
-That is to say, we train a decoder network that generates each cross-correlation waveform $\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}$ using three distinct latent-variable components — they are highlighted in Fig.
-1b.
+That is to say, we train a decoder network that generates each cross-correlation waveform $\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}$ using three distinct latent-variable components — they are highlighted in Figure 1b.
   • Coherent source features $\left(\prescript{b}{\lambda}{\mathbf{g}}_{ij}\right)$.
@@
 $\prescript{b}{}{\widehat{\mathbf{w}}}_{ij}^{(m)}=\mathtt{decoder}\left(\prescript{b}{l}{\mathbf{g}}_{ij},\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}\right),\quad\text{where}\;l=\arg\max_{\lambda}\prescript{b}{\lambda}{\pi}_{ij}^{(m)},\quad m=1,\ldots,M_{ij}.$ (9)
-Here, $l$ denotes the source state with the highest posterior probability for the window $m$, effectively selecting a single coherent feature $\prescript{b}{l}{\mathbf{g}}_{ij}$ for generation.
-We implement the $\mathtt{decoder}$ as a convolutional neural network.
+Here, $l$ denotes the source state with the highest posterior probability for the window $m$.
+The $\mathtt{decoder}$ effectively selects a single coherent feature $\prescript{b}{l}{\mathbf{g}}_{ij}$ for generation of the observed cross-correlation.
+We implement it as a convolutional neural network.
 2.3.2 Inference model
@@
 Here, $\mathtt{enc}_{\mathbf{g}}$, $\mathtt{enc}_{\pi}$, and $\mathtt{enc}_{\mathbf{n}}$ are convolutional encoder components that infer coherent source features, source state probabilities, and nuisance features, respectively.
-The symmetric encoder component $\mathtt{enc}_{\mathbf{g}}$ exploits the permutation invariance by accumulating information across all time windows for a given station pair and branch.
-This component infers $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$, which is constrained to be invariant to window ordering.
+The symmetric encoder component $\mathtt{enc}_{\mathbf{g}}$ is permutation invariant, i.e., its output $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$ is invariant to the ordering of input windows.
+This component accumulates information across all time windows for a given station pair and branch.
 On the contrary, $\mathtt{enc}_{\pi}$ and $\mathtt{enc}_{\mathbf{n}}$ operate on individual windows to capture window-specific features, inferring $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ and $\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}$, respectively.
@@
 After training, we therefore relabeled the source states so that the stationary-zone state is always set to $\lambda=1$.
-Additional architectural and training details are given in the Supplementary Text S22.
+Additional architectural and training details are given in the Supplementary Text S33.
 2.4 Coherent source subsampling (CSS)
@@
 It uses the learned source state probabilities $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ for subsampling.
-Specifically, for each branch $b$, station pair $(i,j)$, and source state $\lambda$, the CSS estimator of the conditional expectation in Eq.
-11 is given by Equation 4, where the subset of correlation windows $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is defined as:
+Specifically, for each branch $b$, station pair $(i,j)$, and source state $\lambda$, the CSS estimator of the conditional expectation in Equation 11 is given by Equation 4, where the subset of correlation windows $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is defined as:
 $\prescript{b}{\lambda}{\mathcal{C}}_{ij}=\left\{m:\prescript{b}{\lambda}{\pi}_{ij}^{(m)}>\alpha\right\}.$ (12)
@@
 With this targeted subsampling and by setting $\lambda=1$, we can isolate the contributions from stationary regions in the source field.
-Note that the conditional average $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ in Eq.
-4 is a non‑linear function of the data, unlike linear averaging, because the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ itself depends on the data.
+Note that the conditional average $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ in Equation 4 is a non‑linear function of the data, unlike linear averaging, because the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ identification itself depends on the data.
 It is important to emphasize that the averaging is performed on the original cross-correlation windows, using the VAE only to identify which windows belong to each source state.
@@
 3 Results
-Figure 2: Cross-correlation gathers for the eastern United States.
-(a) Before subsampling: The linearly averaged cross-correlations, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, show large amplitudes close to zero lag and pronounced causal-acausal asymmetry, indicating a directionally biased ambient noise field.
-(b) After subsampling: The conditional averages, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$, obtained after subsampling, display enhanced waveform coherence and substantially improved causal–acausal symmetry.
-The percentages of time-window allocation (in red) indicate the fraction of windows selected for the source state $\lambda\text{ = 1}$.
-The correlation values (in blue) measure the degree of symmetry between the acausal and causal components for each station pair.
-The black lines mark the time lags corresponding to velocities of $2$ and 55\,km/s.
-[…]
-The improvement in cross-correlation symmetry translates directly into enhanced Rayleigh-wave dispersion measurements.
-Figures 3c–e for (GOGA, NCB) and 3k–m for (BLA, PKME) display the Rayleigh-wave dispersion images obtained without subsampling, which show asymmetry between the causal and acausal branches.
-After applying CSS averaging for $\lambda=1$, the dispersion images (Figures 3f–h and 3n–p) exhibit markedly enhanced coherence and symmetry.
-This improved balance between the causal and acausal branches enables a much more reliable picking of group velocities.
-Figures 4a–i show group-velocity dispersion curves, extracted by automated picking, for nine station pairs with short, intermediate, and long interstation distances.
-For each pair, the dispersion curves obtained by simple linear averaging display substantial scatter and a clear mismatch between the causal and acausal branches.
-After applying conditional averaging, the dispersion estimates become much more stable: the causal and acausal branches closely coincide, producing nearly symmetric behavior in the 5505–50\,s period band, where Rayleigh-wave energy is strongest.
-The pairs with the longest and intermediate distances, including (GOGA, NCB) (14201420\,km), (BLA, PKME) (12901290\,km), (CBN, PKME) (10321032\,km), and (ACSO, CNNC) (713713\,km), show the most pronounced improvements, with the subsampled curves collapsing onto a well-defined trend.
-Other station pairs, such as (AAM, GOGA) (989989\,km), (BINY, GLMI) (755755\,km), (AAM, BINY) (632632\,km), (ACSO, BINY) (625625\,km), and (BLA, GOGA) (505505\,km), also exhibit notably closer agreement between the causal and acausal branches than in the linear averages.
-Finally, we compare the CSS-derived dispersion curves with the earthquake-derived dispersion curves for (GOGA, NCB) and (ACSO, CNNC) pairs, using earthquakes of about $M_{w}\,4.5$ located near the GOGA and ACSO stations, respectively.
-For both pairs, the subsampled symmetric dispersion curves agree well with the earthquake-based dispersion curves.
-It is important to note that while CSS enhances time-domain symmetry, it does not guarantee identical causal and acausal dispersion curves across all frequency bands.
-This is an expected consequence: if the stationary zone lacks sources at specific frequencies, or if certain frequency bands are naturally weak or absent in the ambient field, CSS can enhance the coherence of existing energy but cannot artificially generate it.
-Thus, the method optimizes the extraction of information from the available ambient field but is fundamentally constrained by the spectral content of that field.
-Figure 3: Temporal and dispersion analysis for the station pairs (GOGA, NCB) and (BLA, PKME).
-(a–b, i–j) Acausal and causal time-window selections obtained from CSS for source states $\lambda\text{ = \{1, 2\} }$.
-(c–e, k–m) Before subsampling: standard linear averaging produces dispersion images and cross-correlations with low coherence and pronounced causal–acausal asymmetry.
-(f-h, n-p) After subsampling: source state $\lambda\text{ = 1}$ results in dispersion images and cross-correlations with substantially improved causal–acausal symmetry.
-Figure 4: Group‐velocity dispersion curves for nine station pairs in the eastern United States.
-Solid lines represent the causal and acausal dispersion curves obtained for source state $\lambda\text{ = 1}$, while dotted lines indicate the corresponding dispersion curves based on linear averaging.
-Earthquake‐based dispersion curves are shown for comparison for station pairs (GOGA, NCB) and (ACSO, CNNC).
-For all station pairs, the dispersion curves retrieved using source state $\lambda\text{=1}$ show markedly reduced scatter between 1010\, and 5050\,s, whereas the linearly averaged curves display substantial variability and systematically underestimate velocities.
-This closer agreement demonstrates the effectiveness of CSS for recovering symmetric dispersion information.
+Figure 2: Cross-correlation gathers from the Central California array.
+(a,c) Linear averaging without subsampling, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, for the 3388\,s and 885050\,s bands, respectively.
+The gathers exhibit clear causal–acausal asymmetry, with higher signal-to-noise ratios on the acausal side due to dominant illumination from Pacific Ocean microseisms.
+(b,d) The conditional averaging, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$ using CSS for the same bands, shows improved symmetry between causal and acausal branches.
+Red percentages denote the fraction of time windows assigned to the source state $\lambda=1$ by CSS, and blue values show the correlation between causal and acausal branches for each station pair.
+The black lines mark time lags corresponding to apparent velocities of $1.25$ and 55\,km/s.
+For the Central California array, the correlations after averaging without subsampling, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, shown in Figures 2a and 2c, reveal a strongly directional ambient noise wavefield and display clear causal–acausal asymmetry.
+For both 3388\,s and 885050\,s period bands, the dominant energy is present in the acausal branch, consistent with seismic energy primarily originating from the Pacific Ocean.
+Because the acausal branch exhibits a signal-to-noise ratio (SNR) much higher than that of the causal branch, it is typically used for tomographic imaging.
+However, it should be noted that this measurement may be biased relative to the true interstation response, as indicated by the observed asymmetry.
+In contrast, the conditionally averaged correlations (subsampled averages), $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$, shown in Figures 2b and2d exhibit a markedly stronger symmetry: the Rayleigh waves on the causal and acausal branches are nearly identical.
+This enhanced symmetry is quantified in Figure 2.
+For each pair of stations, Figure 2 also reports the fraction $\frac{\left|\prescript{b}{\lambda}{\mathcal{C}}_{ij}\right|}{M_{ij}}$ of the retained time windows.
+The increased causal–acausal similarity indicates that CSS preferentially selects windows associated with the stationary source illumination state.
+We now obtained correlations that more accurately represent the true interstation response.
+In Figure 3, we show that the enhanced symmetry improves the dispersion measurements.
+Here, the matrices derived without subsampling display pronounced artifacts: the group velocities vary irregularly between adjacent station pairs and across periods, particularly for the low-SNR branch (upper triangles of the matrices in Figures 3a, c and e).
+Although the high-SNR group-velocity picks (lower triangles of Figures 3a, c and e) do not exhibit significant artifacts, they may still be influenced by bias arising from non-uniform source illumination.
+Our subsequent tomography findings will further support the presence of this bias.
+In contrast to linear averaging, we observe that the CSS group-velocity picks are more similar across the branches — note that the artifacts in the low-SNR branch are significantly reduced.
+In addition, the velocity picks in the high-SNR branch change slightly before and after subsampling, which reflects the reduction in bias.
+Consequently, we construct three sets of group‐velocity tomograms (see Figure 4) using: 1.
+picks from the high-SNR branch without subsampling — hereafter, referred to as LS-high tomograms; 2.
+picks obtained without subsampling but after averaging the high- and low-SNR branches — hereafter, referred to as LS tomograms; 3.
+picks derived from the correlations after CSS.
+For CSS tomography, since the group-velocity picks are generally consistent across the different branches, we averaged the values when they matched within a 15% range and discarded the picks otherwise.
+We used the surface‐wave tomography module of SeisLib (Magrini et al., 2022), which performs a regularized least-squares tomography of interstation group‐velocity measurements.
+The regularization parameters were selected to yield comparable apparent smoothness and resolution in all cases so that the differences discussed below arise from the group-velocity picks themselves rather than from inversion settings.
+  • Reduction of structural bias relative to fault geometry: In the LS and LS-high models, velocity anomalies are elongated along the station array and approximately parallel to the dominant raypaths rather than to the mapped faults.
+Such along-array smearing reflects the bias due to non-uniform source distribution.
+In contrast, the CSS tomograms show velocity contrasts that follow the tectonic structure: distinct low-velocity zones adjacent to the RF and the SAF fault system, and a relatively higher-velocity block confined between them.
+This behavior is consistent with controlled-source and local-earthquake imaging studies near Parkfield, which document reduced velocities within the fault damage zone and comparatively higher velocities in adjacent crustal blocks (Hole et al., 2006; Yang et al., 2025) (Supplementary Figure S77).
+The improved alignment of anomalies with mapped structures indicates that CSS reduced source illumination bias and better recovers geologically consistent heterogeneity.
+  • Improved consistency between periods and physical plausibility: Surface-wave group velocities should evolve smoothly with period because sensitivity kernels broaden gradually with depth.
+The CSS models exhibit this expected behavior: the inter-fault block persists from  88\,s to 11~11\,s (roughly 10101313\,km depth) and weakens progressively with increasing period.
+By contrast, LS models display abrupt appearance and disappearance of anomalies and inconsistent velocity changes between neighboring periods.
+Such an inconsistency is unlikely to reflect real structure and instead suggests contamination from period-dependent source distribution, a known limitation of ambient-noise correlation under non-uniform illumination.
+Note that applying smoothing to reduce these inconsistencies will lead to a loss of resolution.
+  • Shallow fault‐zone effects: In shorter periods, CSS tomograms consistently resolve lower group velocities along the RF and SAF fault traces, while the intervening block remains faster.
+This pattern matches the expected seismic signature of a fault damage zone: fractured and fluid-altered rocks reduce seismic velocities within the fault core and surrounding damage region, while more intact crystalline rocks between faults maintain higher velocities.
+Similar low-velocity fault-zone structures have been reported from borehole and seismic studies at SAFOD and related Parkfield experiments (Hole et al., 2006).
+The LS tomograms either smear these regions along the array or fail to isolate them.
+These observations demonstrate that CSS improves ambient noise tomography by recovering structurally coherent and geologically interpretable features.
+We also processed another ambient-noise dataset (Liang and Langston, 2008) from the eastern United States, a stable cratonic region, to provide additional evidence; these results are presented in the Supplementary Text S44.
+Figure 3: Group‐velocity matrices obtained from ambient noise cross-correlations.
+For each subplot, the columns correspond to virtual source stations and rows to receiver stations (CC0101–CC2020).
+The color scale indicates Rayleigh-wave group velocity (km/s).
+Top row (a,c,e) shows picks without subsampling for periods of 88\,s, 99\,s, and 1010\,s, respectively, while the bottom row (b,d,f) shows the corresponding picks after Coherent Source Subsampling (CSS).
+White regions represent the lack of picks e.g., due to a smaller offset relative to Rayleigh-wave wavelength.
+The gray diagonal annotation reports the average similarity between the causal and acausal branch picks.
+Compared to linear averaging, CSS yields significantly higher reciprocity.
+Figure 4: Group‐velocity tomograms (a–l) and ray coverage (m) for the Central California array, where the stations (red inverted triangles) are roughly along the transect marked using the dashed magenta line.
+To obtain the tomograms for periods of 881111\,s (approximate peak sensitivity at depths of  10101313\,km) the group-velocity picks are derived from: the high-SNR branch only (left column), both high- and low-SNR branches (middle column), and the proposed CSS selection (right column).
+Major tectonic elements, including the Southern Coast Ranges (SCR), Rinconada Fault (RF), San Andreas Fault (SAF), and Great Valley Basin (GVB), are indicated.
+The CSS tomograms mirror the underlying geology and reveal a high-velocity block located between the RF and SAF, a feature that is consistently supported by previous studies based on either refraction data or local-earthquake traveltimes.
 4 Discussion
-The results show that CSS effectively isolates the relatively small fraction of ambient-noise windows that exert the strongest control on empirical Green’s function retrieval, yielding substantial improvements in causal–acausal symmetry and dispersion stability.
-By extracting quasi-stable noise regimes that recur over long periods, CSS provides a physically interpretable connection between ambient-noise variability and the theoretical conditions required for reciprocal Green’s function estimation.
-Because it reduces the number of windows needed to achieve coherent stacking, CSS is particularly well suited to large-scale experiments, including dense nodal arrays and distributed acoustic sensing (DAS).
-In such data-heavy environments, CSS can retrieve coherent surface-wave energy even in the presence of intense anthropogenic noise and pronounced spatial heterogeneity, enhancing interferometric quality while greatly decreasing data volume.
-More broadly, the conditional-averaging framework at the core of CSS may benefit applications where coherent wavefield reconstruction is constrained by uncertain or varying source distributions.
-A current practical limitation is that our implementation trains a separate VAE for each station pair; while this is feasible for modest-sized networks, scaling to hundreds of pairs would require substantial computational effort or new joint, multi-pair training strategies.
-Developing such scalable approaches is a crucial step toward deploying CSS at network-wide or continental scales.
-In future work, CSS could also be adapted to independently subsample narrow frequency bands in order to detect frequency-specific weak sources within the stationary zone.
+As CSS can effectively extract interstation response, it is particularly well suited for experiments using dense nodal arrays and distributed acoustic sensing (DAS).
+In these data-intensive settings, the goal of interferometry is to reduce the data volume while preserving useful subsurface information.
+CSS can potentially be used to mitigate biases due to strong directional anthropogenic noise sources in these settings, improving interferometric extraction of interstation response.
+The current practical limitation is that our implementation trains a separate VAE for each station pair; while this is manageable for arrays of moderate size, extending to hundreds of pairs would demand substantial computational resources.
+Development of new, joint multi-pair training schemes is essential as such scalable techniques are a key step toward applying CSS to dense seismic surveys.
 5 Conclusion
-We introduced Coherent Source Subsampling (CSS), a data-driven approach that reformulates ambient noise averaging as a conditional process guided by source-state coherence.
-Instead of relying on indiscriminate stacking, CSS computes conditional averages over subsets representing quasi-stationary segments of the ambient field.
-By isolating these coherent intervals, CSS suppresses the influence of transient and directionally biased noise and substantially enhances the causal–acausal symmetry of the retrieved cross-correlations.
-Application to the eastern United States dataset demonstrates that coherent Green-function estimates can be obtained from only a small fraction of all available time windows.
-Conditional averages produced through CSS yield clearer Rayleigh-wave arrivals, improved causal–acausal agreement, and more stable group-velocity dispersion curves across a range of interstation distances.
-These improvements reduce the bias and variability inherent in linear stacking and provide dispersion measurements better suited for robust ambient-noise tomography.
-In general, CSS offers a physically interpretable and empirically validated pathway for extracting reliable structural information from nonstationary and anisotropic noise fields, extending the applicability of ambient noise imaging in challenging environments.
+We introduced Coherent Source Subsampling (CSS), a data-driven approach that reformulates ambient noise interferometry as a conditional averaging process guided by a neural-network based stationary-zone source selection.
+CSS suppresses the influence of transient and directionally biased noise and substantially enhances the causal–acausal symmetry of the retrieved cross-correlations.
+Application to the central California dataset demonstrates that subsampling yields more reliable Rayleigh-wave arrivals, improved causal–acausal agreement, and more stable group-velocity dispersion curves across a range of interstation distances.
+These improvements reduce the bias and variability inherent to traditional ambient-noise tomography done without subsampling.
+We demonstrated that the CSS tomograms successfully image established geological structures, including a high-velocity block located between the Rinconada and San Andreas faults, as well as low-velocity zones adjacent to these faults.
+In summary, CSS extends the applicability of ambient noise interferometry to challenging environments with complex source distributions and shorter survey durations.
 Open Research Section
 Waveform data used in this study can be obtained from the IRIS EarthScope Data Management Center (https://ds.iris.edu/ds/nodes/dmc/) through the FDSN dataselect service (http://service.iris.edu/fdsnws/dataselect/1/).
-The dataset includes recordings from the United States National Seismic Network (network code US) USArray.
-All codes developed as part of this work will be made publicly available after the manuscript undergoes peer review.
+The dataset includes recordings from the CCSE array (network code TO) (CCSE, 2015) and United States National Seismic Network (network code US) (Albuquerque Seismological Laboratory (ASL)/USGS, 1990).
+The codes developed for this study are included as a ZIP archive in the filetype “Dataset” for peer review and will be made publicly available upon publication.
 Conflict of Interest
@@
 Acknowledgments
-This work was supported by the Shell project under project titled Neural Network Based Denoising Of Passive Seismic Data.
+This work was supported by the Shell project under the project titled Neural Network Based Denoising Of Passive Seismic Data.
 We would like to thank Dr.
@@
 Aurélien Mordret (GEUS), Prof.
-Gerard Schuster (University of Utah) and Dr.
+Gerard Schuster (University of Utah), and Dr.
 Kees Weemstra (KNMI) for their valuable feedback.
-We thank Tiente Rengneichuong Koireng (IISc) and Pushkar Bharadwaj (IISc) for testing the method on other datasets.
+We also thank Prof.
+Brandon Schmandt (Rice University) for discussions on the CCSE dataset.
+We thank Tiente Rengneichuong Koireng (IISc) for testing the method on other datasets.
 The authors used AI-assisted tools (OpenAI ChatGPT and Writefull) for language refinement.